OpenStudy (anonymous):

What is the value of x if e^3x+6=8?

Nnesha (nnesha):

\[\huge\rm e^{3x+6}=8\] how would you cancel out the `e` ? do you know ?

OpenStudy (anonymous):

subtract it on both sides?

OpenStudy (anonymous):

no nvm

Nnesha (nnesha):

no you should take `ln` both sides to cancel out the `e` here is an example \[\huge\rm \cancel{ln e}^x = x\]

Nnesha (nnesha):

if hte question is \[\huge\rm e^9\] i would take ln \[\huge\rm ln e^9\] which cancel out the e answer would by 9

OpenStudy (anonymous):

mc020-2.jpg mc020-3.jpg mc020-4.jpg mc020-5.jpg

Nnesha (nnesha):

???

Nnesha (nnesha):

take ln both sides just like u take square or square root both sides

OpenStudy (anonymous):

i think x= ln6-8/3 ?

Nnesha (nnesha):

ln = natural log

Nnesha (nnesha):

no

OpenStudy (anonymous):

oh x= ln8+6/3

Nnesha (nnesha):

don't post answer choices one by one...

Nnesha (nnesha):

\[\huge\rm lne^{3x+6}=ln8\] noow go from here!

Nnesha (nnesha):

ln and e cancel each other out!

OpenStudy (anonymous):

so 3x+6=ln8 is whats left?

Nnesha (nnesha):

yes right now solve for x!!

Nnesha (nnesha):

how would you cancel out the 6 from the left side ?

OpenStudy (anonymous):

subtract 6

Nnesha (nnesha):

good right so x= ??

OpenStudy (anonymous):

8-6/3

Nnesha (nnesha):

it's not just 8

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\huge\rm lne^{3x+6}=ln8\] noow go from here! \(\color{blue}{\text{End of Quote}}\) ln8 not just 8

OpenStudy (anonymous):

ln8-6/3 srry

Nnesha (nnesha):

looks good!

OpenStudy (anonymous):

thank you